Calculate the size of the multiplier given the following information: (1)
c = 0.4
Government Spending = 100 000
a) 166666.6
b) 1.66
c) 16.6
d) 250 000
1
Expert's answer
2020-10-13T11:06:38-0400

Multiplier = 1/(1-MPC)

Multiplier = 1/(1-0.4) = 1/(0.6) = 1.6666..


Mutiplier is b= 1.66